There are an equal number of males and females in your community, 15% of the population are indigent. At any one time, 45% of the males have high blood pressure. Assume that there is no causal relationship between gender, hypertension, and indigent status.
What is the probability that the next person who comes to your emergency room will be a hypertensive, indigent, man?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

This would be 0.45*0.15=0.0675.

Answer by ikleyn(52778)

.

Tutor @Boreal forgot to account for the gender in his solution.

So the correct answer is P = 0.5*0.15*0.45 = 0.03375.

The first factor/(multiplier)  0.5  in my formula accounts for the man/women possible outcome.